OpenStudy (anonymous):

13 = 5 - 13 +3a (&how did you get the answer ?)

jhonyy9 (jhonyy9):

26-5=3a 21/3=a a=7

OpenStudy (anonymous):

simplify 13=-8+3a add 8 to both sides 21=3a divide by 3 a=7

OpenStudy (anonymous):

13 = -8 +3a 3a = 21 a = 7

OpenStudy (anonymous):

13= 5-13 +3a So,according to the order of operations (PEMDAS), you first figure out the multiplication, but forgetting about it I'll go to 13 - 5, which is 8. 21(3a)-8+13 and that is how you get your answer!!! Therefore, a=7.